# Inhomogeneity, Fluctuations, and Gap Filling in Overdoped Cuprates

### v1

@@ Abstract
Several recent experiments have challenged the premise that cuprate high-temperature superconductors approach conventional Landau-BCS behavior in the high-doping limit. We argue, based on an analysis of their superconducting spectra, that anomalous properties seen in the most-studied overdoped cuprates require a pairing interaction that is strongly inhomogeneous on nm length scales. This is consistent with recent proposals that the “strange-metal” phase above $T_{c}$ in the same doping range arises from a spatially random interaction. We show, via mean-field Bogoliubov-de Gennes (BdG) calculations and time-dependent Ginzburg-Landau (TDGL) simulations, that key features of the observed tunneling spectra are reproduced when both inhomogeneity and thermal phase fluctuations are accounted for. In accord with experiments, BdG calculations find that low-$T$ spectra are highly inhomogeneous and exhibit a low-energy spectral shoulder and broad coherence peaks. However, the spectral gap in this approach becomes homogeneous at high $T$, in contrast to experiments. This is resolved when thermal fluctuations are included; in this case, global phase coherence is lost at the superconducting $T_{c}$ via a broadened BKT transition, while robust phase-coherent superconducting islands persist well above $T_{c}$. The local spectrum remains inhomogeneous at $T_{c}$.
@@
Introduction. Underdoped and optimally doped cuprates represent the epitome of doped Mott physics due to their proximity to the parent insulator, and exhibit not only superconductivity but other intertwined phases as well. Until relatively recently, however, it was generally accepted that the overdoped materials are uncomplicated by a pseudogap or other competing phases and should neatly approach the Landau-BCS paradigm with increasing doping.
@@
Numerous features of the overdoped regime challenge this notion. First, the superfluid density at zero temperature in high quality La2-xSrxCuO4(LSCO) films was found to be proportional to the critical temperature $T_{c}$ [1, 2], in stark contrast to the BCS prediction that it should be equal to the electron density and independent of $T_{c}$. Second, scanning tunneling spectroscopy (STS) experiments in Pb- and La-doped Bi2Sr2CuO6+δ(Bi-2201) repeatedly observe that the spectrum has a two-gap structure, with a homogeneous low-energy gap that is ascribed to superconductivity and an inhomogeneous high-energy gap that is reminiscent of the pseudogap found in underdoped cuprates [3, 4, 5, 6]. A two-gap structure has also been observed in overdoped Bi2Sr2CaCu2O8+δ (Bi-2212) [7, 8], though the two gaps are close in energy and hard to resolve. Third, overdoped Bi-2201 and Bi-2212 are highly inhomogeneous, with a spectral gap that varies by a factor of two or more over distances of a few nm [9, 10, 3, 5, 11, 6]. Temperature-dependent STS [12, 9, 13] and angle-resolved photoemission spectroscopy (ARPES) experiments [14, 15, 16] have shown that in both materials these spectral gaps close well above $T_{c}$, raising questions about the nature of the superconducting state. Magnetoresistance measurements on LSCO found evidence for superconducting fluctuations extending a few tens of kelvin above $T_{c}$ [17] throughout the overdoped regime; from this, it was inferred that $T_{c}$ occurs at a Berezinskii-Kosterlitz-Thouless (BKT) transition. However, magnetic susceptibility measurements [18, 19, 20] in LSCO find diamagnetic signatures extending up to much higher temperatures. These invite a “granular” interpretation, in which superconducting $T_{c}$ is determined by the strength of Josephson coupling between robustly superconducting islands [21].
@@
Furthermore, it has been argued that granularity above $T_{c}$ arises naturally if the overdoped cuprates contain a sufficient density of pair-breaking impurities [20] so that superconductivity persists only in rare regions. However, numerical simulations cast doubt that heavily disordered $d$-wave superconductors are truly granular [22]. Indeed, the disorder levels suggested in Ref. [20] would generate a much larger residual density of states [23] than is actually observed by ARPES measurements on very overdoped Bi-2212 [24], and suppress the average $T_{c}$ to zero.
@@
Figure 1: Plots of $\Delta_{OP}$ (top row) and $\Delta_{SG}$ (middle row) for a single realization of the pairing interaction. The LDOS (bottom row) is binned by spectral gap, and averaged over 50 realizations of the pairing interaction. Results are shown at five temperatures ranging from $T\ll T_{\mathrm{mf}}$ to $T\approx T_{\mathrm{mf}}$, for an impurity concentration of 7.5%. The red arrow indicates the energy scale where the shoulders form.
@@
Here we propose a rather different picture. We suggest that the observed spectroscopic properties of overdoped Bi-2201 and Bi-2212 depend on two ingredients: a localized pairing interaction whose strength varies by an order of magnitude on a scale of nanometers, and inhomogeneous phase fluctuations. We further argue that pair-breaking disorder, which is essential to explain anomalous low-energy properties like the superfluid density [25, 26, 27, 28], is less relevant on the energy scale of the gap. Our calculations show that a two-gap spectrum with a homogeneous subgap and rapidly varying large gap emerges naturally when the inhomogeneity in the pair potential is sufficiently strong, without the need to appeal to a pseudogap, and that both gaps are associated with superconductivity. Furthermore, thermal fluctuations have a dramatic effect on the subgap, which fills with increasing $T$ and disappears when superconducting phase coherence is lost. We find that both Bi-2201 and Bi-2212 can be understood within our model, with the key difference between the two sytems being the degree of inhomogeneity. Our calculations suggest that $T_{c}$ in these cuprates is determined by a BKT-like transition that is broadended by disorder [29, 30] and interlayer coupling [31] rather than a granular transition, and that a weak diamagnetic signal persists above $T_{c}$ because of isolated superconducting islands with anomalously large pairing interactions.
@@
The idea that phase-incoherent pairing can open a partial gap without superconductivity was originally explored in the context of underdoped cuprates [32, 33]. It was deemed an unlikely model for the pseudogap because of the unreasonably large temperature window ($\sim 100$ K) over which a vortex-antivortex plasma must persist [32]. Nonetheless, such models capture coarse features of ARPES and STM data near $T_{c}$ [34, 35, 36, 37, 38, 39, 40, 41, 42]. At low $T$, Nunner et al. [43] and Fang et al. [44] showed that pairing inhomogeneity is responsible for two key properties of spectra in underdoped Bi-2212 [45]—that they are largely particle-hole symmetric and homogeneous at low energies, and that there is an inverse correlation between coherence peak energy and peak height—that cannot be explained by conventional impurity models. The novel feature of the current work is the focus on how large pairing and phase fluctuations modify local spectra in a strongly inhomogeneous superconductor near the transition.
@@
Results: mean-field theory. We solved the self-consistent Bogoliubov-de Gennes (BdG) equations on a $40\times 40$ lattice for a pairing model similar to that of Ref. [43] (see SI for details). The model assumes that the pairing interaction has a weak homogeneous component $V_{0}$ that is modulated locally near randomly-distributed impurities with a Yukawa spatial profile, $V_{I}e^{-r/\lambda}/r$, with $r$ the distance to the impurity and $\lambda$ the range of the modulation. Importantly, $\lambda$ is comparable to the typical superconducting coherence length $\xi$ at $T=0$, so that there are large nanoscale spatial fluctuations of the order parameter.
@@
Figure 1 compares the $d$-wave component of the order parameter, spectral gap, and bin-averaged local density of states (LDOS) for temperatures ranging from near zero to near the mean-field transition temperature, $T_{\mathrm{mf}}$. The order parameter at site $i$ is $\Delta_{OP,i}=|\sum_{\delta}(-1)^{\delta_{y}}\Delta_{i,i+\delta}|$, where $\Delta_{i,i+\delta}$ is the pair wavefunction along the bond connecting sites $i$ and $i+\delta$; the spectral gap $\Delta_{SG,i}$ is the energy of the largest positive-energy peak in the LDOS at site $i$. While $\Delta_{OP,i}$ and $\Delta_{SG,i}$ are shown for a single impurity configuration, the LDOS is obtained from an ensemble of 50 randomly-generated configurations and $15\times 15$ supercells. Each site in the ensemble is assigned a bin based on its value of $\Delta_{SG}$ at $T=0.001$ (energies are given in units of the nearest neighbor hopping, see SI), and the average LDOS is shown for each bin. The same sites are in each bin at all temperatures. Ten bins are used in all; the width of each bin is $0.051$.
@@
Figure 2: Plots of the binned spectral gap (left) and order parameter (right) as functions of temperature. At the lowest measured temperature, each site is sorted its local value of $\Delta_{SG}$ (left) and $\Delta_{OP}$ (right) and put into one of ten bins; the values of $\Delta_{SG}$ and $\Delta_{OP}$ are then averaged over each bin at each temperature. Note that the sites belonging to each bin are different for the two plots.
@@
Approximately 40% of the lattice sites belong to the 1st or 2nd bins. At low $T$, these “small-gap” sites are characterized by a canonical $d$-wave density of states with well-defined coherence peaks. Spectra associated with “large-gap” sites, belonging to the 3rd or higher bin, deviate from a simple $d$-wave form: they have a two-gap structure, with low-energy shoulders and high-energy coherence peaks. The shoulders are often interpreted as a subgap, and we adopt that terminology. The shoulders occur at the same energies for all bins, and coincide with the coherence peak energy of the 2nd bin. We ascribe the shoulders to proximity coupling to small-gap sites; that is, they are the extension of neighboring coherence peaks onto large-gap sites [44]. Where they are distinct, the spectral gap (which coincides with the large gap) reflects the local order parameter, and the subgap reflects the typical order parameter.
@@
The correspondence between $\Delta_{SG}$ and $\Delta_{OP}$ is not perfect: spatial variations of $\Delta_{SG}$ are a factor $\sim 2$ less than $\Delta_{OP}$. Nonetheless, $\Delta_{SG}$ appears granular at high $T$, with robust superconducting islands embedded in a weakly superconducting, nearly uniform background. The granularity is not reflected in the subgap, which is spatially uniform. Indeed, it is striking that an order parameter with spatial variations exceeding an order of magnitude can generate an essentially homogeneous low-energy LDOS. Equally remarkable, the subgap persists to temperatures well above the subgap energy, where superconductivity would be suppressed in a homogeneous BCS superconductor. Its persistence at these elevated temperatures is the result of proximity coupling to nearby regions with large pairing interactions.
@@
Because the typical coherence length is larger than the inter-impurity spacing, this proximity coupling produces a spectrum that is surprisingly homogeneous and $d$-wave-like near $T_{\mathrm{mf}}$. This is evident in the binned LDOS shown in Fig. 1, and is summarized in Fig. 2, which shows the $T$-dependence of both $\Delta_{OP}$ and $\Delta_{SG}$, both sorted into ten equally sized bins according to their respective local values. Both are inhomogeneous at low $T$; however, by $T=0.2\approx T_{\mathrm{mf}}$, $\Delta_{SG}$ is nearly the same for all bins, apart from the largest-gap bin, while $\Delta_{OP}$ varies by an order of magnitude or more between small- and large-gap bins. The homogeneity near $T_{\mathrm{mf}}$ is incompatible with temperature-dependent STS experiments, primarily on Bi-2212 [10, 9], which show that the spectral gap is strongly inhomogeneous near $T_{c}$, and that it fills rather than closes.
@@
Results: fluctuations. To address these discrepancies, we consider order-parameter fluctuations in tandem with inhomogeneity. We first solved a set of time-dependent Ginzburg-Landau (TDGL) equations with thermal noise to obtain an ensemble of order parameter “snapshots.” These calculations naturally allow both amplitude and complex phase fluctuations. Inhomogeneity is built into the model via the lowest-order Ginzburg-Landau (GL) parameter, which has a similar Yukawa spatial dependence as the BdG pairing interaction (see SI). Once order-parameter ensembles were generated, we calculated the LDOS at selected sites using a microscopic non-self-consistent BdG Hamiltonian taking the TDGL snapshots as input. TDGL calculations were for a single disorder realization on a $300\times 300$ lattice. The GL parameters describe a slightly different system than for the BdG calculation; however, the qualitative physics is the same.
@@
Figure 3: Effect of thermal fluctuations on the density of states. Temperature dependence of (a) the root-mean-square correlation function, (b)-(f) the GL local density of states, and (g)-(l) the TDGL local density of states. Insets show the time-averaged phase. In (a), $T_{c}$ is determined from the scaling of the spatial correlation function (see SI). The legend shows the spectral-gap ranges for each bin at $T=0.010$ and the number of sites within each bin.
@@
where $\langle\ldots\rangle_{t}$ indicates a time-average. For the (static) GL calculations, $\overline{\Delta}(T)$ is the spatially-averaged order parameter. In Fig. 3(a), an extrapolation of the low-$T$ slope would suggest a GL transition near $T=0.06t$; however, $\overline{\Delta}(T)$ develops upwards curvature near this temperature (a result of the pairing inhomogeneity; similar behavior is seen in BdG calculations [23]), and only vanishes at $T_{\mathrm{GL}}\approx 0.14t$. For the TDGL calculations, $\overline{\Delta}(T)$ is expected to scale as $L^{-\eta(T)}$ (with $\eta(T)$ a $T$-dependent power and $L$ the linear size of the system) below $T_{c}$ and vanish exponentially above; this provides a simple way to estimate the superconducting transition in finite systems. In Fig. 3(a) we estimate that fluctuations push the transition temperature down by a factor of nearly 3 to $T_{c}\approx 0.05t$. This value is supported by analysis of the the spatial phase correlations (see SI).
@@
The remaining panels in Fig. 3 show the binned LDOS taken from 1000 randomly selected sites, following the same binning scheme as before. At low $T$, the GL spectra are indistinguishable from the TDGL spectra in Fig. 3(g); nearly 75% of the sites belong to the 1st or 2nd bin and have a typical $d$-wave spectrum; these are characterized as “small-gap”. The spectra of the remaining “large-gap” sites have a two-gap structure, with a spatially uniform subgap and an inhomogeneous spectral gap. The subgap and spectral gap are less distinct here than in the BdG calculations because the GL parameters correspond to a less inhomogeneous system. As $T$ is increased [Fig. 3(b)-(f)], the subgap closes, approximately following the temperature dependence of $\overline{\Delta}(T)$.
@@
Spectra from the TDGL calculations are shown in Fig. 3(g)-(l). The magnitude of the time-averaged phase, $|\langle e^{i[\theta_{i}(t)}\rangle_{t}|$, is shown in the insets: it is one (zero) for sites whose correlation time is much longer (shorter) than the simulation time. Thermal fluctuations are increasingly prominent as $T$ increases, and the phase correlations vanish in progressively larger regions of space with increasing $T$. Slightly below $T_{c}$, much of the system is phase-coherent but with patches of phase incoherent pairing, while slightly above $T_{c}$ the pairing is incoherent everywhere except for small islands of phase coherent superconductivity. The effect of these fluctuations is to cause a $T$-dependent filling of the subgap; interestingly, while the binned spectra for the large-gap sites are smoothed by fluctuations, the size of the spectral gap changes very little with increasing temperature. At the highest temperature shown [Fig. 3(i)], there is no trace of superconductivity on $\sim 85\%$ of the sites in the system. Conversely, the sites on which spectra have large gaps, which comprise $\sim 15\%$ of our system, have only partially filled gaps: the proximity-induced subgaps fill quickly with increasing temperature while the high-energy spectral gap persists as a pseudogap [Fig. 3(j)-(l)].

### v2

# Inhomogeneity, Fluctuations, and Gap Filling in Disordered Overdoped Cuprates

@@ Abstract
Several recent experiments have challenged the premise that cuprate high-temperature superconductors approach conventional Landau-BCS behavior in the high-doping limit. We argue, based on an analysis of their superconducting spectra, that anomalous properties seen in the most-studied overdoped cuprates require a pairing interaction that is strongly inhomogeneous on nm length scales. This is consistent with recent proposals that the “strange-metal” phase above $T_{c}$ in the same doping range arises from a spatially random interaction. We show, via mean-field Bogoliubov-de Gennes (BdG) calculations and time-dependent Ginzburg-Landau (TDGL) simulations, that key features of the observed tunneling spectra are reproduced when both inhomogeneity and thermal phase fluctuations are accounted for. In accord with experiments, BdG calculations find that low-$T$ spectra are highly inhomogeneous and exhibit a low-energy spectral shoulder and broad coherence peaks. However, the spectral gap in this approach becomes homogeneous at high $T$, in contrast to experiments. This is resolved when thermal fluctuations are included; in this case, global phase coherence is lost at the superconducting $T_{c}$ via a broadened BKT transition, while robust phase-coherent superconducting islands persist well above $T_{c}$. The local spectrum remains inhomogeneous at $T_{c}$, and the gap is found to fill instead of close with increasing temperature.
I Introduction
@@
Underdoped and optimally doped cuprates represent the epitome of doped Mott physics due to their proximity to the parent insulator, and exhibit not only superconductivity but other intertwined phases as well. Until relatively recently, however, it was generally accepted that the overdoped materials are uncomplicated by a pseudogap or other competing phases and should neatly approach the Landau-BCS paradigm with increasing doping.
@@
Numerous features of the overdoped regime challenge this notion. First, the superfluid density at zero temperature in high quality La2-xSrxCuO4(LSCO) films was found to be proportional to the critical temperature $T_{c}$[Bozovic:2016Dependence, Mahmood:2019microwave], in stark contrast to the BCS prediction that it should be equal to the electron density and independent of $T_{c}$. Second, scanning tunneling spectroscopy (STS) experiments in Pb- and La-doped Bi2Sr2CuO6+δ(Bi-2201) repeatedly observe that the spectrum has a two-gap structure, with a homogeneous low-energy gap that is ascribed to superconductivity and an inhomogeneous high-energy gap that is reminiscent of the pseudogap found in underdoped cuprates [Boyer:2007STM, He:2014STM, Li:2021CDW, Ye:2024Emergent]. A two-gap structure has also been observed in overdoped Bi2Sr2CaCu2O8+δ (Bi-2212) [Alldredge:2008evolution, Pushp:2009STM], though the two gaps are close in energy and hard to resolve. Third, overdoped Bi-2201 and Bi-2212 are highly inhomogeneous, with a spectral gap that varies by a factor of two or more over distances of a few nm [Pasupathy:2008STM, Parker:2010, Boyer:2007STM, Li:2021CDW, Tromp:2023Puddle, Ye:2024Emergent]. Temperature-dependent STS [Gomes2007, Pasupathy:2008STM, Li:2022] and angle-resolved photoemission spectroscopy (ARPES) experiments [hashimoto2014energy, He:2021ARPES, Chen2022] have shown that in both materials these spectral gaps close well above $T_{c}$, raising questions about the nature of the superconducting state. Magnetoresistance measurements on LSCO found evidence for superconducting fluctuations extending a few tens of kelvin above $T_{c}$[Rourke:2011phase] throughout the overdoped regime; from this, it was inferred that $T_{c}$ occurs at a Berezinskii-Kosterlitz-Thouless (BKT) transition. However, magnetic susceptibility measurements [Sonier:2008, Koike:2008inhomogeneous, Li:2021Granular] in LSCO find diamagnetic signatures extending up to much higher temperatures. These invite a “granular” interpretation, in which superconducting $T_{c}$ is determined by the strength of Josephson coupling between robustly superconducting islands [Spivak2009].
@@
Furthermore, it has been argued that granularity above $T_{c}$ arises naturally if the overdoped cuprates contain a sufficient density of pair-breaking impurities [Li:2021Granular] so that superconductivity persists only in rare regions. However, numerical simulations cast doubt that heavily disordered $d$-wave superconductors are truly granular [Breio2022]. Indeed, the disorder levels suggested in Ref. [Li:2021Granular] would generate a much larger residual density of states [Pal:2023] than is actually observed by ARPES measurements on very overdoped Bi-2212 [Yin:2024ARPES], and suppress the average $T_{c}$ to zero.
@@
Figure 1: Spatial maps of the $d$-wave order parameter $\Delta_{OP}$ (top row) and spectral gap $\Delta_{SG}$ (middle row) for a single realization of the pairing interaction. The LDOS (bottom row) is calculated for an ensemble of 50 random pairing realizations and for $15\times 15$ supercells. Sites are binned by the value of $\Delta_{SG}$ at $T=0.001$, and the average LDOS is shown for each bin (bottom row). The same sites are in each bin at all temperatures. Ten equal-width bins spanning $0.085<\Delta_{SG}<0.605$ are used; the legend shows the fraction of sites in each bin and darker colors correspond to lower-energy bins. Results are shown at five temperatures ranging from $T\ll T_{\mathrm{mf}}$ to $T\approx T_{\mathrm{mf}}$, for an impurity concentration of 7.5%. The red arrow in (k) indicates the subgap shoulder.
@@
Here we propose a rather different picture. We suggest that the observed spectroscopic properties of overdoped Bi-2201 and Bi-2212 depend on two ingredients: a localized pairing interaction whose strength varies by an order of magnitude on a scale of nanometers, and inhomogeneous phase fluctuations. We further argue that pair-breaking disorder, which is essential to explain anomalous low-energy properties like the superfluid density [Lee-Hone:2017disorder, Lee-Hone:2018optical, Lee-Hone:2020, Ozdemir:2022], is less relevant on the energy scale of the gap. Our calculations show that a two-gap spectrum with a homogeneous subgap and rapidly varying large gap emerges naturally when the inhomogeneity in the pair potential is sufficiently strong, without the need to appeal to a pseudogap, and that both gaps are associated with superconductivity. Furthermore, thermal fluctuations have a dramatic effect on the subgap, which fills with increasing $T$ and disappears when superconducting phase coherence is lost. We find that both Bi-2201 and Bi-2212 can be understood within our model, with the key difference between the two systems being the degree of inhomogeneity. Our calculations suggest that $T_{c}$ in these cuprates is determined by a BKT-like transition that is broadened by disorder [Benfatto:2009, Mahmood:2022] and interlayer coupling [Benfatto:2007kosterlitz] rather than a granular transition, and that a weak diamagnetic signal persists above $T_{c}$ because of isolated superconducting islands with anomalously large pairing interactions.
@@
The idea that phase-incoherent pairing can open a partial gap without superconductivity was originally explored in the context of underdoped cuprates [franz1998phase, kwon1999effect]. It was deemed an unlikely model for the pseudogap because of the unreasonably large temperature window ($\sim 100$ K) over which a vortex-antivortex plasma must persist [franz1998phase]. Nonetheless, such models capture coarse features of ARPES and scanning tunneling microscopy (STM) data near $T_{c}$[Eckl:2002pair, mayr2005phase, valdez2006single, Chubukov:2007, Berg:2007evolution, valdez2008effects, alvarez2008fermi, Atkinson:2012, Hayward:2014angular]. At low $T$, Nunner et al.[Nunner:2005dopant] and Fang et al.[Fang2006] showed that pairing inhomogeneity is responsible for two key properties of spectra in underdoped Bi-2212 [McElroy:2005]—that they are largely particle-hole symmetric and homogeneous at low energies, and that there is an inverse correlation between coherence peak energy and peak height—that cannot be explained by conventional impurity models. The main feature of the current work is the focus on how large pairing and phase fluctuations modify local spectra in a strongly inhomogeneous superconductor near the transition.
II Results: mean-field theory
We first solved the Bogoliubov-de Gennes (BdG) equations on a $40\times 40$ lattice for the tight-binding Hamiltonian
$H=\sum_{ij\sigma}t_{ij}c^{\dagger}_{i\sigma}c_{j\sigma}+\sum_{ij}\left(\Delta_{ij}^{\ast}c_{j\downarrow}c_{i\uparrow}+\Delta_{ij}c^{\dagger}_{i\uparrow}c^{\dagger}_{j\downarrow}\right),$ (1)
@@ II Results: mean-field theory
with $t_{ij}=t$ ($t_{ij}=-0.35t$) for nearest-neighbor (next-nearest-neighbor) lattice sites $i$ and $j$, and with $\Delta_{ij}=V_{ij}\langle c_{j\downarrow}c_{i\uparrow}\rangle$ determined self-consistently along each bond. $V_{ij}$ is nonzero for nearest-neighbor sites only (see Appendix), and $\Delta_{ij}$ is constrained to be real so that there are no spontaneous supercurrents. We set t=1t=-1 so that $|t|$ is the unit of energy, and the chemical potential is adjusted to maintain an electron density $n=0.85$ per unit cell. Note that the value of $n$ does not matter for this work so long as the Fermi surface is qualitatively reasonable. As in Ref. [Nunner:2005dopant], the model assumes that the pairing interaction $V_{ij}$ has a weak homogeneous component $V_{0}$ that is modulated locally near randomly-distributed impurities with a Yukawa spatial profile, $V_{I}e^{-r/\lambda}/r$, with $r$ the distance to the impurity and $\lambda$ the range of the modulation (see Sec. A.1 for details). Importantly, $\lambda$ is comparable to the typical superconducting coherence length $\xi$ at $T=0$, so that there are large nanoscale spatial fluctuations of the order parameter.
@@
Figure 1 compares the $d$-wave component of the order parameter, spectral gap, and bin-averaged local density of states (LDOS) for temperatures ranging from near zero to near the mean-field transition temperature, $T_{\mathrm{mf}}$. The order parameter at site $i$ is $\Delta_{OP,i}=|\sum_{\delta}(-1)^{\delta_{y}}\Delta_{i,i+\delta}|$, and the spectral gap $\Delta_{SG,i}$ is the energy of the largest positive-energy peak in the LDOS at site $i$. The homogeneous interaction ($V_{0}=0.8$) by itself gives $\Delta_{OP}=0.19$ at $T=0.001$ and has a critical temperature $T_{c}=0.085$.
@@
Figure 2: The binned spectral gap (a) and order parameter (b) as functions of temperature. At the lowest measured temperature, each site is sorted its local value of $\Delta_{SG}$ and put into one of ten bins; the values of $\Delta_{SG}$ and $\Delta_{OP}$ are then averaged over each bin at each temperature.
@@
Approximately 40% of the lattice sites belong to the 1st or 2nd bins, see legend in Fig. 1(o). At low $T$, these “small-gap” sites are characterized by a canonical $d$-wave density of states with well-defined coherence peaks. Spectra associated with “large-gap” sites, belonging to the 3rd or higher bin, deviate from a simple $d$-wave form: they have a two-gap structure, with low-energy shoulders and high-energy coherence peaks. The shoulders are often interpreted as a subgap, and we adopt that terminology. The shoulders occur at the same energies for all bins, and coincide with the coherence peak energy of the 2nd bin. We ascribe the shoulders to proximity coupling to small-gap sites; that is, they are the extension of neighboring coherence peaks onto large-gap sites [Fang2006]. Where they are distinct, the spectral gap (which coincides with the large gap) reflects the local order parameter, and the subgap reflects the typical order parameter.
@@
The correspondence between $\Delta_{SG}$ and $\Delta_{OP}$ is not perfect [Sulangi:2021correlations]: spatial variations of $\Delta_{SG}$ are a factor $\sim 2$ less than $\Delta_{OP}$. Nonetheless, $\Delta_{SG}$ appears granular at high $T$, with robust superconducting islands embedded in a weakly superconducting, nearly uniform background. The granularity is not reflected in the subgap, which is spatially uniform. Indeed, it is striking that an order parameter with spatial variations exceeding an order of magnitude can generate an essentially homogeneous low-energy LDOS. Equally remarkable, the subgap persists to temperatures well above the subgap energy, where superconductivity would be suppressed in a homogeneous BCS superconductor. Its persistence at these elevated temperatures is the result of proximity coupling to nearby regions with large pairing interactions.
@@
Because the typical coherence length is larger than the interimpurity spacing and grows near $T_{c}$, this proximity coupling produces a spectrum that is surprisingly homogeneous and $d$-wave-like near $T_{\mathrm{mf}}$. This is evident in the binned LDOS shown in Fig. 1, and is summarized in Fig. 2, which shows the $T$-dependence of $\Delta_{OP}$ and $\Delta_{SG}$, both sorted into ten equally sized bins according to the local values of $\Delta_{SG}$. Both are inhomogeneous at low $T$; however, by $T=0.2\approx T_{\mathrm{mf}}$, $\Delta_{SG}$ is nearly the same for all bins, apart from the largest-gap bin, while $\Delta_{OP}$ varies by an order of magnitude or more between small- and large-gap bins. The homogeneity of $\Delta_{SG}$ near $T_{\mathrm{mf}}$ is incompatible with temperature-dependent STS experiments, primarily on Bi-2212 [Parker:2010, Pasupathy:2008STM], which show that the spectral gap is strongly inhomogeneous near $T_{c}$, and that it fills rather than closes.
III Results: fluctuations
@@
To address these discrepancies, we consider order-parameter fluctuations in tandem with inhomogeneity. We first solved a set of time-dependent Ginzburg-Landau (TDGL) equations with thermal noise to obtain an ensemble of order parameter “snapshots.” These calculations naturally allow both amplitude and complex phase fluctuations. Inhomogeneity is built into the model via the lowest-order Ginzburg-Landau (GL) parameter, which has a similar Yukawa spatial dependence as the BdG pairing interaction (see section A.1). Once order-parameter ensembles were generated, we calculated the LDOS at selected sites using a microscopic non-self-consistent BdG Hamiltonian, with the same tight-binding parameters as above and taking the TDGL snapshots as input. TDGL calculations were for a single disorder realization on a $300\times 300$ lattice.
This approach follows a broad body of work in which an ansatz is made to generate an ensemble of order parameters that capture the essential ingredients of pair fluctuations across the phase transition. Most of the published work (see e.g. Refs. [franz1998phase, kwon1999effect, Eckl:2002pair, mayr2005phase, valdez2006single, Chubukov:2007, Berg:2007evolution, valdez2008effects, alvarez2008fermi, Atkinson:2012, Hayward:2014angular]) takes their ensembles from classical Monte Carlo simulations of a GL free energy or two-dimensional XY model. The TDGL calculations are closely related to classical Monte Carlo approaches in that they both generate pair fluctuations that are consistent with the equipartition theorem and they both generate vortex-antivortex excitations near $T_{c}$. The explicit time in the TDGL calculations is important for conductivity calculations since it determines the frequency-dependent response to an applied EM field, but is not relevant for the density of states since tunneling is an essentially instantaneous process. For our purposes, there is no difference between the real TDGL time and fictitious “Monte Carlo time.” The two approaches contain the same physics, and we chose the TDGL approach for computational convenience.
Superconducting fluctuations contain a longitudinal component that does not satisfy the time-independent equation $\nabla\cdot{\bf J}=0$, but should satisfy the full continuity equation $\partial_{t}\rho+\nabla\cdot{\bf J}=0$. In all of the phenomenological approaches mentioned above, the microscopic BdG Hamiltonian violates the continuity equation for any given order-parameter snapshot, but satisfies it in the ensemble average. It is widely believed that this phenomenological approach accurately captures the essential qualitative physics of the superconducting spectrum. Indeed, the physics is intuitive: currents generated by phase fluctuations shift the quasiparticle excitation spectrum up or down in energy (so-called Doppler shifts), which broadens features in the density of states and produces gap-filling. (See Franz and Millis [franz1998phase] for a discussion of this physics.)
@@
Figure 3: Effect of thermal fluctuations on the density of states. Temperature dependence of (a) the root-mean-square correlation function, (b)-(f) the GL local density of states, and (g)-(l) the TDGL local density of states. Insets show the time-averaged phase. In (a), $T_{c}$ is determined from the scaling of the spatial correlation function (see Section C.2). The legend shows the spectral-gap ranges for each bin at $T=0.010$ and the number of sites within each bin.
@@
where $\langle\ldots\rangle_{t}$ indicates a time-average. For the (static) GL calculations, $\overline{\Delta}(T)$ is the spatially-averaged order parameter. In Fig. 3(a), an extrapolation of the low-$T$ slope would suggest a GL transition near $T=0.06t$; however, $\overline{\Delta}(T)$ develops upwards curvature near this temperature (a result of the pairing inhomogeneity; similar behavior is seen in BdG calculations [Pal:2023]), and only vanishes at $T_{\mathrm{GL}}\approx 0.14t$. For the TDGL calculations, $\overline{\Delta}(T)$ is expected to scale as $L^{-\eta(T)}$ (with $\eta(T)$ a $T$-dependent power and $L$ the linear size of the system) below $T_{c}$ and vanish exponentially above; this provides a simple way to estimate the superconducting transition in finite systems. In Fig. 3(a) we estimate that fluctuations push the transition temperature down by a factor of nearly 3 to $T_{c}\approx 0.05t$. This value is supported by analysis of the the spatial phase correlations (see Fig. 4).
@@
The remaining panels in Fig. 3 show the binned LDOS taken from 1000 randomly selected sites, following the same binning scheme as before. At low $T$, the GL spectra are indistinguishable from the TDGL spectra in Fig. 3(g); $\sim 50\%$ of the sites belong to the 1st or 2nd bin and have a typical $d$-wave spectrum; these are characterized as “small-gap”. The spectra of the remaining “large-gap” sites have a two-gap structure, with a spatially uniform subgap and an inhomogeneous spectral gap. The subgap and spectral gap are less distinct here than in the BdG calculations because the GL parameters correspond to a less inhomogeneous system. As $T$ is increased [Fig. 3(b)-(f)], the subgap closes, approximately following the temperature dependence of $\overline{\Delta}(T)$.
@@
Spectra from the TDGL calculations are shown in Fig. 3(g)-(l). The magnitude of the time-averaged phase, $|\langle e^{i\theta_{i}(t)}\rangle_{t}|$, is shown in the insets: it is one (zero) for sites whose correlation time is much longer (shorter) than the simulation time. Thermal fluctuations are increasingly prominent as $T$ increases, and the phase correlations vanish in progressively larger regions of space with increasing $T$. Slightly below $T_{c}$, much of the system is phase-coherent but with patches of phase incoherent pairing, while slightly above $T_{c}$ the pairing is incoherent everywhere except for small islands of phase coherent superconductivity. The effect of these fluctuations is to cause a $T$-dependent filling of the subgap; interestingly, while the binned spectra for the large-gap sites are smoothed by fluctuations, the size of the spectral gap changes very little with increasing temperature. At the highest temperature shown [Fig. 3(i)], there is no trace of superconductivity on $\sim 75\%$ of the sites in the system. Conversely, the sites on which spectra have large gaps, which comprise $\sim 25\%$ of our system, have only partially filled gaps: the proximity-induced subgaps fill quickly with increasing temperature while the high-energy spectral gap persists as a pseudogap [Fig. 3(j)-(l)].
IV Discussion